How much money will need to invest today if it will be compounded bi-monthly with the rate of 6 % and will be withdrawn in 3 years? The amount of money after three years is ₱35425.



Compute the total interest of a money worth ₱2500 after 7 months if is based on simple interest with the rate of 11%.



Find the amount of ₱13450 after 5 years if it will be considered compounded semi annually at the rate of 9%.


1
Expert's answer
2022-03-16T10:16:23-0400

PRINCIPAL

Let the amount to invest (P) = P

Let the accrued amount (A) = 35425

The annual rate of interest (R) = 0.06

Let the compounded times (n) = 6

Let time in years (t) =3

"A= P(1+r\/n)"nt

"35425= P(1+0.06\/6)"(6*3)

"35425= P(1+0.01)" 18

"35425= P(1.196147476)"

Therefore "P=(35425\/1.196147476)"

P = 29615.91336

The amount required for investment in order to get an accrued amount of 35425 from a compound interest at a rate of 6% over three years is approximately 29615.91.



TOTAL INTEREST

Let the amount invested (P) = 2500

Interest rate (R) =11%

Time (T) = 7 months

Simple Interest = "P*\t\u200b\n \\frac{R}{100}*T"

"= 2500*\t\u200b\n \\frac{11}{100}* \\frac{7}{12}"

= 160.4166667

Total simple interest is approximately 160.42.


ACCRUED AMOUNT

Principal (P)= 13450

Interest rate (R) = 9%

compounded times (n) = 10

Time (t) = 5

Let time in years (t) =3

"A= P(1+r\/n)"nt

"A= 13450(1+0.09\/2)"(2*5)

"A= 13450(1+0.045)" 10

"A= 13450(1.552969422)"

Therefore "A=20887.43872"

The total accrued amount with a principal of 13450 with an interest rate 0f 9% per year compounded semi annually over 5 years will be approximately 20887.44.